# ConsistCompose: Unified Multimodal Layout Control for Image Composition

### v2

@@ 1 Introduction
Recent advances in diffusion models [ho2020denoising] and large-scale vision–language training have substantially improved the quality and versatility of image generation. Systems such as DALL-E [ramesh2021zero] and Stable Diffusion [rombach2022high] demonstrate impressive photorealism, while recent unified multimodal models [xie2024show, zhou2024transfusion, wu2025janus, zhang2025unified, pan2025transfer, chen2025janus, deng2025emerging] integrate perception, reasoning, and generation into a single architecture. Despite their broad capabilities, these unified models remain predominantly focused on visual understanding—especially grounding [fu2025llmdet, wang2025learning, jiang2025detect]—and provide limited support for generating images that satisfy explicit spatial constraints, which is a prerequisite for controllable and compositional scene synthesis.
Layout-controlled generation has emerged as an important direction for enabling spatially grounded image synthesis. However, existing approaches present fundamental obstacles to unification. Diffusion-based methods typically rely on layout–image fusion modules or region-aware U-Net modifications [li2023gligen, wang2024instancediffusion, Zhou2024MIGCMG, zhou2024migc++, wang2025msdiffusion, peng2025muse], making their architectures tightly coupled to layout-specific designs and difficult to reconcile with Transformer-based generative frameworks. Multimodal or autoregressive models such as LayoutSAM [zhang2025creatilayout], HiCo [bocheng2024hico], SMARLI [zheng2025layout], and Plangen [he2025plangen] attempt to incorporate layout either as an independent modality or through structured spatial tokens, and some explore unified backbones for layout planning and layout-to-image generation. Nevertheless, these systems remain restricted to layout-centric tasks and have not demonstrated compatibility with broader multimodal abilities such as visual reasoning [shao2024visual], image understanding, reference grounding, or editing. Moreover, most prior work focuses solely on text-conditioned layout control and seldom considers the more challenging setting of multi-reference, identity-preserving multi-instance composition. This limitation is further compounded by the lack of large-scale datasets that jointly provide instance-level layouts, subject correspondences, and diverse multi-instance configurations, limiting progress toward truly unified layout-aware multimodal systems.
@@
Our experiments show that ConsistCompose delivers state-of-the-art layout controllability, achieving a 7.2% gain in layout IoU and a 13.7% AP improvement on COCO-Position. Beyond COCO-Position, it exhibits strong identity preservation under multi-reference layouts on MS-Bench [wang2025msdiffusion], while maintaining MMMU and MMBench performance on par with its Bagel backbone. These results confirm that explicit spatial control can be incorporated into unified multimodal generation without sacrificing general capabilities.
@@ 2 Related Work
Text-to-Image Generation. Text-to-image (T2I) generation has progressed rapidly with the development of diffusion models [ho2020denoising, podell2023sdxl, ramesh2021zero, rombach2022high, esser2024scaling, cao2025hunyuanimage, Wu2025QwenImageTR], enabling high-fidelity and semantically coherent image generation across diverse domains. However, purely text-driven guidance offers limited control over spatial composition and structural layout, restricting its applicability in scenarios that require precise object placement, poses, or scene configurations.
@@
Layout-Controlled Image Generation. Beyond dense-structure conditioning methods such as ControlNet [zhang2023adding], T2I-Adapter [mou2024t2i], and related approaches [voynov2023sketch, tan2025ominicontrol, xiao2025omnigen], which enforce geometric priors via edges or segmentation maps, layout-based approaches [li2023gligen, wang2024instancediffusion, bocheng2024hico, zhang2025creatilayout, Zhou2024MIGCMG, zhou2024migc++] offer a more flexible means of spatial control. GLIGEN [li2023gligen] introduces bounding-box constraints through gated Transformer layers for open-set generalization, while InstanceDiffusion [wang2024instancediffusion] achieves instance-level control via multimodal input fusion. MIGC [Zhou2024MIGCMG, zhou2024migc++] mitigates attribute leakage and spatial misalignment with a divide-and-conquer attention mechanism, and CreatiLayout [zhang2025creatilayout] treats layout as an independent modality, employing SiamLayout to balance text–layout interactions for improved consistency. Recent studies further extend this line toward multi-instance generation guided by both layout and visual references: MUSE [peng2025muse] fuses layout and instance information via concatenated Cross-Attention; MS-Diffusion [wang2025msdiffusion] introduces a Grounding Resampler with Multi-Subject Cross-Attention for coherent multi-subject synthesis; and ContextGen [xu2025contextgen] enhances spatial–appearance alignment through Contextual Layout Anchoring and Identity Consistency Attention. Despite their progress, these approaches remain fragmented and highly task-specific, often relying on auxiliary fusion modules or handcrafted attention mechanisms. Such designs limit scalability and hinder integration into unified multimodal frameworks. In contrast, our work explores LELG under a unified multimodal paradigm, embedding layout semantics directly into language prompts to achieve consistent, layout-aware, and identity-preserving multi-instance generation.
@@
Unified Multimodal Generation. The recent rise of multimodal large language models (MLLMs) has enabled unified architectures that perform both understanding and generation within a single framework [zhang2025unified]. In parallel, diffusion-based models have evolved toward unified multimodal generation, where a single model supports diverse generation tasks under shared objectives [tan2025ominicontrol, tan2025ominicontrol2, xiao2025omnigen, wu2025omnigen2explorationadvancedmultimodal]. The release of GPT-4o [hurst2024gpt] further underscores the potential of such cross-modal unification. Early efforts, including Chameleon [team2024chameleon] and Emu2 [Emu2], extend language modeling to multimodal contexts by treating images as token sequences and applying next-token prediction across modalities. Emu3 [wang2024emu3] enhances this design via discrete VQ-VAE tokenization, enabling unified autoregressive modeling within a shared token space. Hybrid frameworks such as Show-O [xie2024show] and Transfusion [zhou2024transfusion] combine Transformers and diffusion within a joint pipeline to integrate token prediction and denoising. Janus [wu2025janus] and Janus Pro [chen2025janus] reveal that a single visual encoder struggles to balance perception and generation, while MetaQuery [pan2025transfer] and BLIP-3o [chen2025blip3] encode conditional information via fixed query tokens, which limits representational capacity and leads to information loss for complex prompts. OmniGen2 [wu2025omnigen2explorationadvancedmultimodal] alleviates this issue by directly leveraging interleaved multimodal hidden states from MLLMs as diffusion inputs, avoiding token-level compression.
@@
Inspired by the MMDiT mechanism, recent systems such as Bagel [deng2025emerging] and Mogao [liao2025mogao] adopt a deep-fusion architecture that jointly processes textual and visual tokens within shared self-attention layers, achieving fine-grained modality interaction and improved semantic alignment. Building upon this evolution, we further investigate the MoT architecture and extend its deep-fusion mechanism for controllable multi-instance image generation. Our framework targets the key challenge of layout-controlled multimodal generation—ensuring spatial precision, compositional consistency, and identity preservation—through LELG, which embeds layout semantics and instance identities directly within the unified vision-language interface under a single generative paradigm.
@@ 3 Method
We present ConsistCompose, a unified multimodal generation framework built upon the MoT architecture of Bagel [deng2025emerging]. At its core lies the proposed LELG paradigm, which enables spatially controllable multi-instance generation by embedding explicit layout semantics—encoded as textual tokens—directly into the linguistic stream. This design allows layout control to be natively integrated into a unified multimodal modeling pipeline without introducing task-specific or layout-centric branches. An overview of the framework is shown in Fig. 1.
@@ 3.1 Model Architecture
ConsistCompose adopts Bagel as its backbone, featuring two Transformer experts within the MoT architecture: one for multimodal understanding and the other for multimodal generation. Two visual encoders are employed: a SigLIP-initialized ViT [tschannen2025siglip] for semantic perception and a FLUX-initialized VAE [labs2025flux] for image generation. Both experts operate on a shared token sequence through unified self-attention, ensuring consistent latent representations across perception and generation. All modalities—including text, ViT features, and VAE latents—are projected into a common embedding space for seamless multimodal fusion:
@@ 3.2 Linguistic-Embedded Layout-Grounded Generation
This text-embedded spatial representation constitutes the Instance-Coordinate Binding Prompt (ICBP), a minimal yet effective mechanism that binds each object’s linguistic reference to its spatial coordinates within a unified generative sequence.
@@ 3.4 Training Objectives
Training combines Flow Matching for image generation and autoregressive token prediction for multimodal understanding. Given a text-image pair $(\mathbf{c},\mathbf{x})$, Flow Matching [lipman2022flow] interpolates between latent states:
@@
We follow the standard formulation of Flow Matching [lipman2022flow] without architectural modifications.
@@ 3.6 ConsistCompose3M Dataset
The T2I component is obtained by reprocessing LayoutSAM [zhang2025creatilayout], which provides diverse multi-object layouts with instance-level annotations. For each instance, we append its normalized bounding box to the caption via the ICBP mechanism (Sec. 3.2), injecting layout coordinates into the prompt and turning dense layout annotations into language-embedded supervision without extra coordinate channels or layout-specific encoders. The reference-conditioned split targets identity-preserving layout control. We reuse subject-centric assets from existing pipelines—subject banks and crops from Subjects200K [tan2025ominicontrol] and UNO [wu2025less], defined on the Objects365 taxonomy—rather than generating new subjects. These subjects are recomposed into multi-subject scenes under diverse layouts, with bounding boxes and explicit pairings between each scene and its reference images. We then attach ICBP-style captions and filter samples with CLIP/DINO-based similarity to enforce identity consistency and discard layout-inconsistent or low-quality examples. Taken together, these two components constitute a large-scale corpus with tightly aligned spatial and semantic supervision, supporting both text-only and reference-conditioned layout-aware generation and providing a principled testbed for studying layout-controllable generation in unified multimodal models.
4 Experiment
@@
Methods Instance Success Ratio $(\%)\uparrow$ Image Success Ratio $(\%)\uparrow$ Position Accuracy $(\%)\uparrow$
@@
Annotation GLIGEN InstanceDiffusion MIGC MIGC++ CreatiLayout (FLUX) Alignment Hybrid SFT
@@
Figure 4: Qualitative Comparison on COCO-Position: ConsistCompose Respects Target Layouts with Coherent Detailed Scenes.
We conduct extensive experiments to evaluate ConsistCompose on both layout-controlled generation and general multimodal understanding. We benchmark on COCO-Position [lin2014microsoft, Zhou2024MIGCMG] and MS-Bench [wang2025msdiffusion] for text- and vision-conditioned multi-instance layout control, and assess general capabilities on GenEval [ghosh2023geneval], MMMU [yue2024mmmu], MMBench [liu2024mmbench], and DreamBench [ruiz2023dreambooth]. Together, these evaluations test (i) fine-grained spatial control, (ii) identity-preserving multi-subject generation, and (iii) retention of broad reasoning ability under unified multimodal training.
@@ 4.1 Implementation Details
Training Data. We initialize our model from the Bagel foundation model and primarily train it on ConsistCompose3M. To preserve broad visual–linguistic understanding, we additionally incorporate FineVision [Wiedmann2025FineVisionOD] and MAmmoTH [yue2023mammoth]. To better support multi-instance layout reasoning, we further mine high-resolution (>512>512 px) images from COCO2017 [lin2015microsoftcococommonobjects] and Objects365 [shao2019objects365], discarding low-quality samples. These corpora provide dense multi-object scenes and same-category groupings that strengthen the model’s understanding of parallel layouts within the LELG framework.
In the second-stage hybrid SFT, we interleave two complementary datasets: GPT-Image-Edit-1.5M [wang2025gptimageedit15mmillionscalegptgeneratedimage] to retain editing capabilities and Text-to-Image-2M [jackyhate_text-to-image-2M_2024] to further enhance text-driven generation. This two-stage schedule allows us to inject layout-awareness while preserving the strong general abilities of the base model.
4.2 Evaluation Metrics and Benchmarks
COCO-Position. Spatial accuracy on COCO-Position is evaluated using Instance Success Ratio, Image Success Ratio, and Position Accuracy. Instance Success Ratio is computed over instances whose predicted bounding boxes achieve $\text{IoU}>0.5$ (via Grounding DINO [liu2024grounding]) and is reported across different instance levels $L_{2}$$L_{6}$. Image Success Ratio additionally requires all instances in an image to be successfully localized, capturing global layout coherence. Position Accuracy is measured by mIoU and AP, reflecting fine-grained alignment between predicted and target box locations across varying IoU thresholds.
MS-Bench. MS-Bench [peng2025muse, wang2025msdiffusion] contains 1,148 multi-subject combinations (2-3 subjects per image) from 40 categories, each with ground-truth bounding boxes. We follow prior work and evaluate three aspects: (1) positional accuracy, (2) text-image consistency, and (3) identity consistency. CLIP-T [hessel2022clipscorereferencefreeevaluationmetric] measures global text-image semantic alignment, DINOv2 [oquab2023dinov2] captures fine-grained correspondence between cropped generations and reference subjects, and IoU/AP (via Grounding DINO) quantify localization accuracy under multi-subject layouts. We report results on both the original MS-Bench split and the MS-Bench-Random variant, which shuffles object layouts to stress-test spatial generalization while keeping subject identities fixed.
@@ 4.2 Evaluation Metrics and Benchmarks
General Capabilities. To verify that layout-aware training does not degrade broad multimodal competence, we evaluate on MMMU, MMBench, GenEval, and DreamBench [yue2024mmmu, liu2024mmbench, ghosh2023geneval, ruiz2023dreambooth]. MMMU is reported with Task Accuracy to reflect multi-domain reasoning; MMBench uses the Overall Comprehension Score to evaluate visual understanding and cross-modal grounding; GenEval provides object-centric metrics for counting, spatial relations, attribute binding, and negation. DreamBench assesses subject-driven generation quality using CLIP-T (global semantics), CLIP-I (local visual-semantic matching), and DINO (fine-grained feature similarity) under single- and multi-object settings.

### v3

@@ Abstract
Figure 1: Key Results of ConsistCompose: Advancing Layout-Controllable Multi-Instance Generation. (a) Layout-Grounded Text-to-Image Generation with Faithful Prompt Alignment and High Precision. (b) Content-Preserving, Flexible Rearrangement of Image Elements. (c) Multi-Instance Layout Control for Structured Complex Scene Generation with Identity Preservation.
@@ 1 Introduction
Recent advances in diffusion models [13] and large-scale vision–language training have substantially improved the quality and versatility of image generation. Systems such as DALL-E [31] and Stable Diffusion [32] demonstrate impressive photorealism, while recent unified multimodal models [53, 61, 49, 59, 28, 3, 7] integrate perception, reasoning, and generation into a single architecture. Despite their broad capabilities, these unified models remain predominantly focused on visual understanding—especially grounding [9, 42, 16]—and provide limited support for generating images that satisfy explicit spatial constraints, which is a prerequisite for controllable and compositional scene synthesis.
Layout-controlled generation has emerged as an important direction for enabling spatially grounded image synthesis. However, existing approaches present fundamental obstacles to unification. Diffusion-based methods typically rely on layout–image fusion modules or region-aware U-Net modifications [18, 45, 62, 63, 43, 29], making their architectures tightly coupled to layout-specific designs and difficult to reconcile with Transformer-based generative frameworks. Multimodal or autoregressive models such as LayoutSAM [57], HiCo [4], SMARLI [60], and PlanGen [11] attempt to incorporate layout either as an independent modality or through structured spatial tokens, and some explore unified backbones for layout planning and layout-to-image generation. Nevertheless, these systems remain restricted to layout-centric tasks and have not demonstrated compatibility with broader multimodal abilities such as visual reasoning [34], image understanding, reference grounding, or editing. Moreover, most prior work focuses solely on text-conditioned layout control and seldom considers the more challenging setting of multi-reference, identity-preserving multi-instance composition. This limitation is further compounded by the lack of large-scale datasets that jointly provide instance-level layouts, subject correspondences, and diverse multi-instance configurations, limiting progress toward truly unified layout-aware multimodal systems.
@@
Our experiments show that ConsistCompose delivers state-of-the-art layout controllability, achieving a 7.2% gain in layout IoU and a 13.7% AP improvement on COCO-Position. Beyond COCO-Position, it exhibits strong identity preservation under multi-reference layouts on MS-Bench [43], while maintaining MMMU and MMBench performance on par with its Bagel backbone. These results confirm that explicit spatial control can be incorporated into unified multimodal generation without sacrificing general capabilities.
@@ 2 Related Work
Text-to-Image Generation. Text-to-image (T2I) generation has progressed rapidly with the development of diffusion models [13, 30, 31, 32, 8, 1, 48], enabling high-fidelity and semantically coherent image generation across diverse domains. However, purely text-driven guidance offers limited control over spatial composition and structural layout, restricting its applicability in scenarios that require precise object placement, poses, or scene configurations.
@@
Layout-Controlled Image Generation. Beyond dense-structure conditioning methods such as ControlNet [58], T2I-Adapter [26], and related approaches [41, 37, 52], which enforce geometric priors via edges or segmentation maps, layout-based approaches [18, 45, 4, 57, 62, 63] offer a more flexible means of spatial control. GLIGEN [18] introduces bounding-box constraints through gated Transformer layers for open-set generalization, while InstanceDiffusion [45] achieves instance-level control via multimodal input fusion. MIGC [62, 63] mitigates attribute leakage and spatial misalignment with a divide-and-conquer attention mechanism, and CreatiLayout [57] treats layout as an independent modality, employing SiamLayout to balance text–layout interactions for improved consistency. Recent studies further extend this line toward multi-instance generation guided by both layout and visual references: MUSE [29] fuses layout and instance information via concatenated Cross-Attention; MS-Diffusion [43] introduces a Grounding Resampler with Multi-Subject Cross-Attention for coherent multi-subject synthesis; and ContextGen [54] enhances spatial–appearance alignment through Contextual Layout Anchoring and Identity Consistency Attention. Despite their progress, these approaches remain fragmented and highly task-specific, often relying on auxiliary fusion modules or handcrafted attention mechanisms. Such designs limit scalability and hinder integration into unified multimodal frameworks. In contrast, our work explores LELG under a unified multimodal paradigm, embedding layout semantics directly into language prompts to achieve consistent, layout-aware, and identity-preserving multi-instance generation.
@@
Unified Multimodal Generation. The recent rise of multimodal large language models (MLLMs) has enabled unified architectures that perform both understanding and generation within a single framework [59]. In parallel, diffusion-based models have evolved toward unified multimodal generation, where a single model supports diverse generation tasks under shared objectives [37, 38, 52, 50]. The release of GPT-4o [14] further underscores the potential of such cross-modal unification. Early efforts, including Chameleon [39] and Emu2 [36], extend language modeling to multimodal contexts by treating images as token sequences and applying next-token prediction across modalities. Emu3 [44] enhances this design via discrete VQ-VAE tokenization, enabling unified autoregressive modeling within a shared token space. Hybrid frameworks such as Show-O [53] and Transfusion [61] combine Transformers and diffusion within a joint pipeline to integrate token prediction and denoising. Janus [49] and Janus Pro [3] reveal that a single visual encoder struggles to balance perception and generation, while MetaQuery [28] and BLIP-3o [2] encode conditional information via fixed query tokens, which limits representational capacity and leads to information loss for complex prompts. OmniGen2 [50] alleviates this issue by directly leveraging interleaved multimodal hidden states from MLLMs as diffusion inputs, avoiding token-level compression.
@@
Inspired by the MMDiT mechanism, recent systems such as Bagel [7] and Mogao [19] adopt a deep-fusion architecture that jointly processes textual and visual tokens within shared self-attention layers, achieving fine-grained modality interaction and improved semantic alignment. Building upon this evolution, we further investigate the MoT architecture and extend its deep-fusion mechanism for controllable multi-instance image generation. Our framework targets the key challenge of layout-controlled multimodal generation—ensuring spatial precision, compositional consistency, and identity preservation—through LELG, which embeds layout semantics and instance identities directly within the unified vision-language interface under a single generative paradigm.
@@ 3 Method
We present ConsistCompose, a unified multimodal generation framework built upon the MoT architecture of Bagel [7]. At its core lies the proposed LELG paradigm, which enables spatially controllable multi-instance generation by embedding explicit layout semantics—encoded as textual tokens—directly into the linguistic stream. This design allows layout control to be natively integrated into a unified multimodal modeling pipeline without introducing task-specific or layout-centric branches. An overview of the framework is shown in Fig. 2.
@@ 3.1 Model Architecture
ConsistCompose adopts Bagel as its backbone, featuring two Transformer experts within the MoT architecture: one for multimodal understanding and the other for multimodal generation. Two visual encoders are employed: a SigLIP-initialized ViT [40] for semantic perception and a FLUX-initialized VAE [17] for image generation. Both experts operate on a shared token sequence through unified self-attention, ensuring consistent latent representations across perception and generation. All modalities—including text, ViT features, and VAE latents—are projected into a common embedding space for seamless multimodal fusion:
@@ 3.2 Linguistic-Embedded Layout-Grounded Generation
This text-embedded spatial representation forms the Instance Coordinate Binding Prompt (ICBP), a minimal yet effective mechanism that binds each object’s linguistic reference to its spatial coordinates within a unified generative sequence.
@@ 3.4 Training Objectives
Training combines Flow Matching for image generation and autoregressive token prediction for multimodal understanding. Given a text-image pair $(\mathbf{c},\mathbf{x})$, Flow Matching [22] interpolates between latent states:
@@
We follow the standard formulation of Flow Matching [22] without architectural modifications.
@@ 3.6 ConsistCompose3M Dataset
Methods Instance Success Ratio (%) $\uparrow$ Image Success Ratio (%) $\uparrow$ Position Accuracy (%) $\uparrow$
@@
Annotation GLIGEN InstanceDiffusion MIGC++ CreatiLayout PlanGen Alignment Hybrid SFT
@@
Figure 5: Qualitative Comparison on COCO-Position: ConsistCompose respects target layouts with coherent and detailed scenes.
The T2I component is obtained by reprocessing LayoutSAM [57], which provides diverse multi-object layouts with instance-level annotations. For each instance, we append its normalized bounding box to the caption via the ICBP mechanism (Sec. 3.2), injecting layout coordinates into the prompt and turning dense layout annotations into language-embedded supervision without extra coordinate channels or layout-specific encoders. The reference-conditioned split targets identity-preserving layout control. We reuse subject-centric assets from existing pipelines—subject banks and crops from Subjects200K [37] and UNO [51], defined on the Objects365 taxonomy—rather than generating new subjects. These subjects are recomposed into multi-subject scenes under diverse layouts, with bounding boxes and explicit pairings between each scene and its reference images. We then attach ICBP-style captions and filter samples with CLIP/DINO-based similarity to enforce identity consistency and discard layout-inconsistent or low-quality examples. Taken together, these two components constitute a large-scale corpus with tightly aligned spatial and semantic supervision, supporting both text-only and reference-conditioned layout-aware generation and providing a principled testbed for studying layout-controllable generation in unified multimodal models.
4 Experiment
We conduct extensive experiments to evaluate ConsistCompose on both layout-controlled generation and general multimodal understanding. We benchmark on COCO-Position [21, 62] and MS-Bench [43] for text- and vision-conditioned multi-instance layout control, and assess general capabilities on GenEval [10], GEdit [24], MMMU [55], MMBench [25], and DreamBench [33]. Together, these evaluations test (i) fine-grained spatial control, (ii) identity-preserving multi-subject generation, and (iii) retention of broad reasoning ability under unified multimodal training.
@@ 4.1 Implementation Details
Training Data. We initialize our model from the Bagel foundation model and primarily train it on ConsistCompose3M. To preserve broad visual–linguistic understanding, we additionally incorporate FineVision [47] and MAmmoTH [56]. To better support multi-instance layout reasoning, we further mine high-resolution (>512>512 px) images from COCO2017 [20] and Objects365 [35], discarding low-quality samples. These corpora provide dense multi-object scenes and same-category groupings that strengthen the model’s understanding of parallel layouts within the LELG framework.
In the second-stage hybrid SFT, we interleave two complementary datasets: GPT-Image-Edit-1.5M [46] to retain editing capabilities and Text-to-Image-2M [15] to further enhance text-driven generation. This two-stage schedule allows us to inject layout-awareness while preserving the strong general abilities of the base model.
4.2 Evaluation Metrics and Benchmarks
COCO-Position. Spatial accuracy on COCO-Position is evaluated using Instance Success Ratio, Image Success Ratio, and Position Accuracy. Instance Success Ratio is computed over instances whose predicted bounding boxes achieve $\text{IoU}>0.5$ (via Grounding DINO [23]) and is reported across different instance levels $L_{2}$$L_{6}$. Image Success Ratio additionally requires all instances in an image to be successfully localized, capturing global layout coherence. Position Accuracy is measured by mIoU and AP, reflecting fine-grained alignment between predicted and target box locations across varying IoU thresholds.
MS-Bench. MS-Bench [29, 43] contains 1,148 multi-subject combinations (2-3 subjects per image) from 40 categories, each with ground-truth bounding boxes. We follow prior work and evaluate three aspects: (1) positional accuracy, (2) text-image consistency, and (3) identity consistency. CLIP-T [12] measures global text-image semantic alignment, DINOv2 [27] captures fine-grained correspondence between cropped generations and reference subjects, and IoU/AP (via Grounding DINO) quantify localization accuracy under multi-subject layouts. We report results on both the original MS-Bench split and the MS-Bench-Random variant, which shuffles object layouts to stress-test spatial generalization while keeping subject identities fixed.
@@ 4.2 Evaluation Metrics and Benchmarks
General Capabilities. To verify that layout-aware training does not degrade broad multimodal competence, we evaluate on MMMU, MMBench, GenEval, GEdit, and DreamBench [55, 25, 10, 24, 33]. MMMU is reported with Task Accuracy to reflect multi-domain reasoning; MMBench uses the Overall Comprehension Score to evaluate visual understanding and cross-modal grounding; GenEval provides object-centric metrics for counting, spatial relations, attribute binding, and negation; GEdit evaluates instruction-driven image editing quality; DreamBench assesses subject-driven generation under single- and multi-object settings.